Calculate the mole fraction of solute in the aqueous solution of 13% C2H5OH.


1
Expert's answer
2022-05-26T10:10:04-0400

The easiest way in case of percent is to start with a mass of 100 gram solution. This solution contains 13 g ethanol and 87 g water.


13 g ethanol = 13/ 46 = 0.28 moles ethanol


87 g water = 87 / 18 = 4.83 moles water


the mole fraction = 0.28 / (0.28 + 4.83) = 0.05